What is the value of x when: 2x + 10 = 6 - 9x
Is it:
2x + 9x = 6 - 10
11x = -4, x = -4/11
Is that right? I am so confused ^-_-^

Answer by rothauserc(4718)

2x + 10 = 6 - 9x
add 9x to both sides of the =
11x +10 = 6
subtract 10 from both sides of the =
11x = -4
divide both sides of = by 11
x = -4/11
**************
check the answer by substituting for x in the given equation
2*(-4/11) +10 = 6 - 9*(-4/11)
-8/11 + 10 = 6 + 36/11
convert whole numbers to fractions
-8/11 + 110/11 = 66/11 + 36/11
102/11 = 102/11
our answer checks